# Quasi-Bayes empirical Bayes: a sequential approach to the Poisson compound decision problem

### v1

@@ D.1 Synthetic data
We set $n=50,100,200,500$, and generate $(Y_{1},\ldots,Y_{n})$ under the Poisson compound model, with $(\theta_{1},\ldots,\theta_{n})$ i.i.d from a Uniform prior $G$ on $[0,3]$. For the NP-QB method, we set Newton’s algorithm as follows: i) the parameter $\theta$ takes value in $\Theta=\{0.01,0.02,\ldots,99.99,100\}$; ii) the number $d$ of mixture components is the cardinality of $\Theta$, i.e. $d=10,000$; iii) the initial guess $g_{0}$ is the Uniform distribution on $\Theta$; iv) the learning rate is $\alpha_{n}=(1+n)^{-1}$. Results are displayed in Figure 2 and Table 3.
@@
Figure 2: Data $Y_{i}$’s against the $\theta_{i}$’s (black circle), and estimates of the $\theta_{i}$’s: NP-EB (green line), NP-ML (cyan line), NP-MD (magenta line), P-EB (red line) and NP-QB (blue line).
Table 3: Values of RMSE and MAD for the NP-EB, NP-ML, NP-MD, P-EB and NP-QB estimates.
NP-EB NP-ML NP-MD P-EB NP-QB
RMSE   $n=100$ 1.01 0.75 1.31 0.75 0.75
RMSE   $n=500$ 0.95 0.75 0.74 0.75 0.78
MAD     $n=100$ 0.77 0.60 1.05 0.61 0.60
MAD     $n=500$ 0.65 0.59 0.59 0.60 0.61
@@
Figure 3: Data $Y_{i}$’s against the $\theta_{i}$’s (black circle), and estimates of the $\theta_{i}$’s: NP-EB (green line), NP-ML (cyan line), NP-MD (magenta line), P-EB (red line) and NP-QB (blue line).
Table 4: Values of RMSE and MAD for the NP-EB, NP-ML, NP-MD, P-EB and NP-QB estimates.
NP-EB NP-ML NP-MD P-EB NP-QB
RMSE   $n=100$ 5.83 1.99 1.71 1.83 1.75
RMSE   $n=500$ 2.63 1.89 1.86 1.93 1.86
MAD     $n=100$ 3.40 1.52 1.32 1.38 1.34
MAD     $n=500$ 1.75 1.42 1.42 1.47 1.43
D.2 Real data: Tweeter dataset
We consider a dataset of tweets with at least 50 retweets on Twitter in the period October 7 to November 6, 2011 (30 days). The dataset is available at the link https://snap.stanford.edu/seismic/, and it consists of $166,783$ tweets that generate $34,617,704$ retweets. More specifically, the data provides information on the tweet ID, the tweet post time, the retweet post time and the number of followers of the poster/retweeter.
@@ D.2 Real data: Tweeter dataset
For each tweet, we focus on the number of retweets generated within $30$ seconds from the tweet post time. From the first two rows of Table 5, $40,259$ tweets have 00 retweets in $30$ seconds, $28,339$ tweets have $1$ retweet in $30$ seconds, etc. Table 5 and Figure 4 display the estimates of the number of retweets expected for a tweet.
@@ E An extension to the multidimensional Poisson compound decision problem
The next result extends Theorem 2.2 to the multidimensional case.
@@
We now consider the Bayesian regret
@@
under a “true” Poisson compound model for $(\mathbf{Y}_{n})_{n\geq 1}$, under which, for any $n\geq 1$: i) the distribution of $\mathbf{Y}_{n}$ given $\bm{\theta}^{\ast}_{n}$ is a product of Poisson distributions with parameters $\theta^{\ast}_{n,j}$; ii) the $\bm{\theta}^{\ast}_{n}$’s are independent and identically distributed from the “true” prior distribution, say the oracle prior $G^{\ast}$ on $\Theta_{d}^{k}$, with probability mass function $g^{\ast}$ on $\Theta_{d}^{k}$. The asymptotic behavior of regret for large $n$ in the multidimensional case follows the same rate of convergence as in the univariate case.
@@
Let $g_{n}$ be as in (E.10) with $\alpha_{n}=(\alpha+n)^{-\gamma}$ for $\alpha>0$ and $\gamma\in(1/2,1)$, and let $g^{\ast}$ be the probability mass function of the oracle prior $G^{\ast}$. Denote by $\mathds{P}^{*}$ a probability measure under which the “true” Poisson compound model for $(\mathbf{Y}_{n})_{n\geq 1}$ holds true. Then, for every $\delta<2-1/\gamma$, it holds that
$\text{Regret}(g_{n},g^{*})=o(n^{-\delta})\quad\mathds{P}^{\ast}\,\text{-}\,a.s.$
@@ F.1 Notation
and initial state $\mathbf{g}_{0}\in\mathring{(\Delta^{D})}$, converges $\mathds{P}^{*}$-a.s. to $\mathbf{g}^{*}$ as $n\rightarrow\infty$.
@@
Let $J=\partial\Delta^{D}\cup\{\mathbf{g}^{*}\}$. We first prove that ${\rm d}(\mathbf{g}_{n},J)\rightarrow 0$, $\mathds{P}^{*}$-a.s., as $n\rightarrow\infty$, where ${\rm d}$ denotes the Euclidean distance in $\mathbb{R}^{D-1}$. The result follows from Theorem 2.2.1′′ in Chen (2005) if we can prove that
  • A1
@@
  • A3
    $\lim_{T\rightarrow\infty}\limsup_{n\rightarrow\infty}\frac{1}{T}||\sum_{i=n}^{m(n,T)}\alpha_{i+1}{\bm{\epsilon}}_{i+1}||=0,$ $\mathds{P}^{*}$-a.s., where $m(n,T)=\max\{m:\sum_{i=n}^{m}\alpha_{i+1}\leq T\}$,
  • A4
@@
A1 holds by assumption; A3 holds since $({\bm{\epsilon}}_{n})$ is a martingale difference, and A4 is obvious. We now prove A2. The set $J$ is closed and $\mathscr{l}$ is continuously differentiable. Moreover, for $\mathbf{g}\in\mathring{(\Delta^{D})}$,
@@
which leads to a contradiction. Hence A2 holds, and ${\rm d}(\mathbf{g}_{n},J)\rightarrow 0$ $\mathds{P}^{*}$-a.s., as $n\rightarrow\infty$.
To conclude, we need to show that $\mathds{P}^{*}(\liminf_{n}{\rm d}(\mathbf{g}_{n},\partial\Delta^{D})=0)=0$. Since ${\rm d}(\mathbf{g}^{*},\partial\Delta^{D})>0$ and ${\rm d}(\mathbf{g}_{n},\{\mathbf{g}^{*}\}\cup\partial\Delta^{D})\rightarrow 0$ $\mathds{P}-a.s.$, then $\mathds{P}^{*}(\liminf_{n}{\rm d}(\mathbf{g}_{n},\partial\Delta^{D})=0)>0$ if and only if there exists $M$ with $\mathds{P}^{*}(M)>0$ such that ${\rm d}(\mathbf{g}_{n}(\omega),\partial\Delta^{D})\rightarrow 0$ for every $\omega\in M$, as $n\rightarrow\infty$. Let us make this assumption as a contradiction hypothesis. Define the function $\mathcal{L}(\mathbf{g})=\sum_{i=1}^{D}\log(\mathbf{g}^{*}_{i}/\mathbf{g}_{i})\mathbf{g}^{*}_{i}$. By the contradiction hypothesis, $\mathcal{L}(\mathbf{g}_{n})(\omega)\rightarrow\infty$ as $n\rightarrow\infty$, for every $\omega\in M$. Then $E^{*}({\mathcal{L}}(\mathbf{g}_{n}))\rightarrow\infty$ as $n\rightarrow+\infty$, where $E^{*}$ denote expectation with respect to $\mathds{P}^{*}$. Let $\tilde{\mathbf{g}}_{n}=\eta_{n}\mathbf{g}_{n}+(1-\eta_{n})\mathbf{g}_{n+1}$, with the random variable $\eta_{n}$ taking values in $[0,1]$, be such that
@@
$E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\right)=\alpha_{n+1}E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{h}_{i}(\mathbf{g}_{n})+{\bm{\epsilon}}_{n+1,i})\right),$
@@
$E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}_{i}}{\mathbf{g}_{n,i}}{\bm{\epsilon}}_{n+1,i}\right)=0,$

### v2

@@ D.1 Synthetic data
Table 3: Weibull prior: grid $\Theta_{n,\eta}$ as $n$ and $\eta$ vary.
$\eta=0.025$ $\eta=0.05$ $\eta=0.075$ $\eta=0.1$
$n=50$ $d_{\eta}=454,169$ $d_{\eta}=104,053$ $d_{\eta}=43,755$ $d_{\eta}=23,612$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=11,354.225$ $\vartheta_{d_{\eta}}=52,02.65$ $\vartheta_{d_{\eta}}=3,281.625$ $\vartheta_{d_{\eta}}=2,361.2$
$n=100$ $d_{\eta}=369,148$ $d_{\eta}=84,375$ $d_{\eta}=35,423$ $d_{\eta}=19,090$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=9,228.7$ $\vartheta_{d_{\eta}}=4,218.75$ $\vartheta_{d_{\eta}}=2,656.725$ $\vartheta_{d_{\eta}}=1,909$
$n=200$ $d_{\eta}=450,573$ $d_{\eta}=103,220$ $d_{\eta}=43,402$ $d_{\eta}=23,420$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=11,264.325$ $\vartheta_{d_{\eta}}=5,161$ $\vartheta_{d_{\eta}}=3,255.15$ $\vartheta_{d_{\eta}}=2,342$
$n=500$ $d_{\eta}=431,768$ $d_{\eta}=98,865$ $d_{\eta}=41,558$ $d_{\eta}=22,419$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=10,794.2$ $\vartheta_{d_{\eta}}=49,43.25$ $\vartheta_{d_{\eta}}=31,16.85$ $\vartheta_{d_{\eta}}=2,241.9$
Table 4: Weibull prior: RMSE and MAD on the grid $\Theta_{n,\eta}$ as $n$ and $\eta$ vary.
$\eta=0.025$ $\eta=0.05$ $\eta=0.075$ $\eta=0.1$
$n=50$
RMSE 1.572 1.575 1.578 1.581
MAD 1.319 1.322 1.325 1.328
$n=100$
RMSE 1.677 1.72 1.75 1.772
MAD 1.335 1.366 1.386 1.4
$n=200$
RMSE 1.335 1.336 1.336 1.337
MAD 1.064 1.064 1.065 1.065
$n=500$
RMSE 1.348 1.348 1.349 1.349
MAD 1.071 1.071 1.071 1.071
[…]
Table 5: Weibull prior: RMSE, MAD and CPU time (in seconds) on the grid $\Theta_{n,d}$ as $n$ and $d$ vary.
$d=50,000$ $d=10,000$ $d=5,000$ $d=1,000$ $d=500$
$n=50$
RMSE 1.574 1.581 1.582 6.13 14.49
MAD 1.323 1.328 1.329 5.706 12.518
CPU time 1.243 0.24 0.12 0.025 0.013
$n=100$
RMSE 1.667 1.675 1.677 4.496 11.476
MAD 1.335 1.335 1.335 4.101 10.536
CPU time 1.208 0.24 0.12 0.024 0.013
$n=200$
RMSE 1.33 1.336 1.336 6.039 13.984
MAD 1.062 1.064 1.065 5.733 11.949
CPU time 1.231 0.242 0.122 0.025 0.013
$n=500$
RMSE 1.348 1.348 1.348 5.869 12.96
MAD 1.071 1.071 1.072 5.553 11.152
CPU time 1.214 0.241 0.12 0.025 0.013
Table 6 compares the RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB. The QB-EB implements a Newton’s algorithm is initialized as follows: i) the parameter $\theta$ takes values on a grid $\Theta_{n,d}=\{\vartheta_{1},\ldots,\vartheta_{d}\}$, with $\vartheta_{1}$ and $\vartheta_{d}$ being the minimum and maximum value of $\Theta_{n,0.025}$ in Table 3, respectively, and $d=10,000$; ii) the initial guess $g_{0}$ is the Uniform distribution on $\Theta_{n,d}$; iii) the learning rate is $\alpha_{n}=(1+n)^{-0.99}$. This is precisely the initialization in the second column of Table 5, which also provides CPU time.
Table 6: Weibull prior: RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB.
NP-EB NP-ML NP-MD P-EB QB-EB
RMSE   $n=50$ 3.154 1.542 1.559 1.554 1.581
RMSE   $n=100$ 2.516 1.695 1.648 1.478 1.675
RMSE   $n=200$ 3.236 1.273 1.248 1.262 1.336
RMSE   $n=500$ 2.108 1.341 1.515 1.353 1.348
MAD     $n=50$ 2.32 1.262 1.258 1.295 1.328
MAD     $n=100$ 1.807 1.334 1.3 1.179 1.335
MAD     $n=200$ 1.917 0.999 1.001 1.014 1.064
MAD     $n=500$ 1.4 1.068 1.197 1.073 1.071
[…]
Table 7: Uniform prior: grid $\Theta_{n,\eta}$ as $n$ and $\eta$ vary.
$\eta=0.025$ $\eta=0.05$ $\eta=0.075$ $\eta=0.1$
$n=50$ $d_{\eta}=71,674$ $d_{\eta}=15,988$ $d_{\eta}=6,596$ $d_{\eta}=3,504$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=1,791.85$ $\vartheta_{d_{\eta}}=799.4$ $\vartheta_{d_{\eta}}=494.7$ $\vartheta_{d_{\eta}}=350.4$
$n=100$ $d_{\eta}=67,264$ $d_{\eta}=14,986$ $d_{\eta}=6,177$ $d_{\eta}=3,279$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=1,681.6$ $\vartheta_{d_{\eta}}=749.3$ $\vartheta_{d_{\eta}}=463.275$ $\vartheta_{d_{\eta}}=327.9$
$n=200$ $d_{\eta}=52,298$ $d_{\eta}=11,593$ $d_{\eta}=4,760$ $d_{\eta}=2,519$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=1,307.45$ $\vartheta_{d_{\eta}}=579.65$ $\vartheta_{d_{\eta}}=357$ $\vartheta_{d_{\eta}}=251.9$
$n=500$ $d_{\eta}=44,591$ $d_{\eta}=9,850$ $d_{\eta}=4,034$ $d_{\eta}=2,130$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=1,114.775$ $\vartheta_{d_{\eta}}=492.5$ $\vartheta_{d_{\eta}}=302.55$ $\vartheta_{d_{\eta}}=213$
Table 8: Uniform prior: RMSE and MAD on the grid $\Theta_{n,\eta}$ as $n$ and $\eta$ vary.
$\eta=0.025$ $\eta=0.05$ $\eta=0.075$ $\eta=0.1$
$n=50$
RMSE 0.724 0.725 0.727 0.728
MAD 0.615 0.616 0.617 0.618
$n=100$
RMSE 0.798 0.798 0.799 0.8
MAD 0.648 0.65 0.651 0.653
$n=200$
RMSE 0.728 0.729 0.73 0.731
MAD 0.619 0.62 0.621 0.622
$n=500$
RMSE 0.72 0.72 0.721 0.721
MAD 0.59 0.59 0.59 0.59
[…]
Table 9: Uniform prior: RMSE, MAD and CPU time (in seconds) on the grid $\Theta_{n,d}$ as $n$ and $d$ vary.
$d=50,000$ $d=10,000$ $d=5,000$ $d=1,000$ $d=500$
$n=50$
RMSE 0.724 0.724 0.724 0.737 1.587
MAD 0.615 0.615 0.616 0.619 1.397
CPU time 1.233 0.241 0.125 0.026 0.014
$n=100$
RMSE 0.799 0.804 0.812 0.841 1.514
MAD 0.649 0.653 0.66 0.678 1.302
CPU time 1.310 0.241 0.120 0.026 0.014
$n=200$
RMSE 0.728 0.73 0.731 0.751 1.056
MAD 0.619 0.619 0.619 0.623 0.831
CPU time 1.252 0.239 0.131 0.025 0.013
$n=500$
RMSE 0.72 0.719 0.719 0.721 0.854
MAD 0.59 0.612 0.612 0.64 0.665
CPU time 1.207 0.241 0.119 0.025 0.013
Table 10 compares the RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB. The QB-EB implements a Newton’s algorithm is initialized as follows: i) the parameter $\theta$ takes values on a grid $\Theta_{n,d}=\{\vartheta_{1},\ldots,\vartheta_{d}\}$, with $\vartheta_{1}$ and $\vartheta_{d}$ being the minimum and maximum value of $\Theta_{n,0.025}$ in Table 7, respectively, and $d=10,000$; ii) the initial guess $g_{0}$ is the Uniform distribution on $\Theta_{n,d}$; iii) the learning rate is $\alpha_{n}=(1+n)^{-0.99}$. This is precisely the initialization considered in the second column of Table 9, which also provides the CPU time. Figure 2 displays the NP-EB, NP-ML, NP-MD, P-EB and QB-EB estimates.
Table 10: Uniform prior: RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB.
NP-EB NP-ML NP-MD P-EB QB-EB
RMSE   $n=50$ 1.354 0.727 0.702 0.703 0.724
RMSE   $n=100$ 1.253 0.818 0.841 0.763 0.804
RMSE   $n=200$ 0.88 0.718 0.729 0.726 0.73
RMSE   $n=500$ 0.88 0.725 0.746 0.72 0.719
MAD     $n=50$ 0.996 0.608 0.562 0.581 0.615
MAD     $n=100$ 0.998 0.673 0.685 0.628 0.653
MAD     $n=200$ 0.742 0.613 0.605 0.623 0.619
MAD     $n=500$ 0.719 0.606 0.633 0.607 0.612
@@
Figure 2: Data $Y_{i}$’s (Uniform prior) against the $\theta_{i}$’s (black circle), and estimates of the $\theta_{i}$’s: NP-EB (green line), NP-ML (cyan line), NP-MD (magenta line), P-EB (red line) and QB-EB (blue line).
[…]
Table 11: Half-Gaussian prior: grid $\Theta_{\eta}$
$\eta=0.025$ $\eta=0.05$ $\eta=0.075$ $\eta=0.1$
$n=50$ $d_{\eta}=29,100$ $d_{\eta}=6,362$ $d_{\eta}=2,585$ $d_{\eta}=1,355$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=727.5$ $\vartheta_{d_{\eta}}=318.1$ $\vartheta_{d_{\eta}}=193.875$ $\vartheta_{d_{\eta}}=135.5$
$n=100$ $d_{\eta}=19,629$ $d_{\eta}=4,243$ $d_{\eta}=1,709$ $d_{\eta}=889$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=490.725$ $\vartheta_{d_{\eta}}=212.15$ $\vartheta_{d_{\eta}}=128.175$ $\vartheta_{d_{\eta}}=88.9$
$n=200$ $d_{\eta}=16,878$ $d_{\eta}=3,631$ $d_{\eta}=1,456$ $d_{\eta}=755$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=421.95$ $\vartheta_{d_{\eta}}=181.55$ $\vartheta_{d_{\eta}}=109.2$ $\vartheta_{d_{\eta}}=75.5$
$n=500$ $d_{\eta}=17,307$ $d_{\eta}=3,727$ $d_{\eta}=1,496$ $d_{\eta}=776$
$\vartheta_{1}=0.025$ $\vartheta_{1}=0.05$ $\vartheta_{1}=0.075$ $\vartheta_{1}=0.1$
$\vartheta_{d_{\eta}}=432.675$ $\vartheta_{d_{\eta}}=186.35$ $\vartheta_{d_{\eta}}=112.2$ $\vartheta_{d_{\eta}}=77.6$
Table 12: Half-Gaussian prior: RMSE and MAD on the grid $\Theta_{n,\eta}$ as $n$ and $\eta$ vary.
$\eta=0.025$ $\eta=0.05$ $\eta=0.075$ $\eta=0.1$
$n=50$
RMSE 0.569 0.57 0.571 0.572
MAD 0.462 0.463 0.464 0.465
$n=100$
RMSE 0.633 0.633 0.627 0.624
MAD 0.532 0.535 0.538 0.541
$n=200$
RMSE 0.471 0.471 0.472 0.472
MAD 0.372 0.373 0.374 0.375
$n=500$
RMSE 0.536 0.537 0.537 0.538
MAD 0.425 0.425 0.426 0.427
[…]
Table 13: Half-Gaussian prior: RMSE, MAD and CPU time (in seconds) on the grid $\Theta_{n,d}$ as $n$ and $d$ vary.
$d=50,000$ $d=10,000$ $d=5,000$ $d=1,000$ $d=500$
$n=50$
RMSE 0.569 0.57 0.57 0.58 0.632
MAD 0.461 0.462 0.462 0.463 0.489
CPU time 1.234 0.242 0.134 0.027 0.013
$n=100$
RMSE 0.633 0.633 0.633 0.635 0.687
MAD 0.541 0.541 0.542 0.543 0.594
CPU time 1.226 0.241 0.12 0.024 0.013
$n=200$
RMSE 0.471 0.471 0.471 0.476 0.486
MAD 0.372 0.372 0.372 0.375 0.385
CPU time 1.208 0.241 0.118 0.024 0.013
$n=500$
RMSE 0.535 0.535 0.535 0.536 0.536
MAD 0.419 0.42 0.424 0.424 0.425
CPU time 1.223 0.241 0.12 0.025 0.013
Table 14 compares the RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB. The QB-EB implements a Newton’s algorithm is initialized as follows: i) the parameter $\theta$ takes values on a grid $\Theta_{n,d}=\{\vartheta_{1},\ldots,\vartheta_{d}\}$, with $\vartheta_{1}$ and $\vartheta_{d}$ being the minimum and maximum value of $\Theta_{n,0.025}$ in Table 11, respectively, and $d=10,000$; ii) the initial guess $g_{0}$ is the Uniform distribution on $\Theta_{n,d}$; iii) the learning rate is $\alpha_{n}=(1+n)^{-0.99}$. This is precisely the initialization considered in the second column of Table 13, which also provides the CPU time. Figure 2 displays the NP-EB, NP-ML, NP-MD, P-EB and QB-EB estimates.
Table 14: Half-Gaussian prior: RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB.
NP-EB NP-ML NP-MD P-EB QB-EB
RMSE   $n=50$ 1.129 0.623 0.643 0.578 0.57
RMSE   $n=100$ 0.679 0.511 0.484 0.493 0.633
RMSE   $n=200$ 0.611 0.477 0.491 0.483 0.471
RMSE   $n=500$ 0.838 0.555 0.564 0.54 0.535
MAD     $n=50$ 0.773 0.49 0.493 0.467 0.462
MAD     $n=100$ 0.512 0.428 0.391 0.411 0.541
MAD     $n=200$ 0.467 0.374 0.403 0.388 0.372
MAD     $n=500$ 0.508 0.429 0.44 0.43 0.42
@@
Figure 3: ata $Y_{i}$’s (half-Gaussian prior) against the $\theta_{i}$’s (black circle), and estimates of the $\theta_{i}$’s: NP-EB (green line), NP-ML (cyan line), NP-MD (magenta line), P-EB (red line) and QB-EB (blue line).
D.2 Real data: Twitter dataset
We analyze a dataset of tweets that received at least $50$ retweets on Twitter between October 7 and November 6, 2011, available at https://snap.stanford.edu/seismic/. It contains $166,783$ tweets generating $34,617,704$ retweets, along with metadata such as tweet ID, posting time, retweet time, and follower counts. For each tweet, we focus on the number of retweets received within the first $30$ seconds after posting.
@@ D.2 Real data: Twitter dataset
For each tweet, we consider the number of retweets generated within $30$ seconds from the tweet post time. From the first two rows of Table 15, $40,259$ tweets have 00 retweets in $30$ seconds, $28,339$ tweets have $1$ retweet in $30$ seconds, etc. Table 15 and Figure 4 display the estimates of the number of retweets expected for a tweet.
@@ E An extension to the multidimensional Poisson compound decision problem
The next result extends Theorem 2.3 in two directions: the observations are now random vectors, and the sequence of learning rates $(\alpha_{n})$ is no longer required to be increasing.
@@
We now consider the Bayesian regret under a “true” Poisson compound model for $(\mathbf{Y}_{n})_{n\geq 1}$, under which, for any $n\geq 1$: i) the distribution of $\mathbf{Y}_{n}$ given $\bm{\theta}^{\ast}_{n}$ is a product of Poisson distributions with parameters $\theta^{\ast}_{n,j}$; ii) the $\bm{\theta}^{\ast}_{n}$’s are independent and identically distributed from the “true” prior distribution, say the oracle prior $G^{\ast}$. By the same techniques used in the univariate case, the frequentist guarantees can be studied in this more general setting. We give a complete treatment under the simplifying assumption that the “true” probability distribution $G^{*}$ has support $\Theta_{d}^{k}$. Denoting by $g^{*}$ its probability mass function on $\Theta_{d}^{k}$, the regret becomes
@@
The asymptotic behavior of regret for large $n$ in the multidimensional case follows the same rate of convergence as in the univariate case.
@@
Let $g_{n}$ be as in (E.19) with $\alpha_{n}=(\alpha+n)^{-\gamma}$ for $\alpha>0$ and $\gamma\in(1/2,1)$, and let $g^{\ast}$ be the probability mass function of the oracle prior $G^{\ast}$. Denote by $\mathds{P}_{G^{*}}$ a probability measure under which the “true” Poisson compound model for $(\mathbf{Y}_{n})_{n\geq 1}$ holds true. Then, for every $\delta<2-1/\gamma$, it holds that
$\text{Regret}(g_{n},g^{*})=o(n^{-\delta})\quad\mathds{P}_{G^{*}}\,\text{-}\,a.s.$
@@ F.1 Notation
and initial state $\mathbf{g}_{0}\in\mathring{(\Delta^{D})}$, converges $\mathds{P}_{G^{*}}$-a.s. to $\mathbf{g}^{*}$ as $n\rightarrow\infty$.
@@
Let $J=\partial\Delta^{D}\cup\{\mathbf{g}^{*}\}$. We first prove that ${\rm d}(\mathbf{g}_{n},J)\rightarrow 0$, $\mathds{P}_{G^{*}}$-a.s., as $n\rightarrow\infty$, where ${\rm d}$ denotes the Euclidean distance in $\mathbb{R}^{D-1}$. The result follows from Theorem 2.2.1′′ in Chen (2005) if we can prove that
  • S1
@@
  • S3
    $\lim_{T\rightarrow\infty}\limsup_{n\rightarrow\infty}\frac{1}{T}||\sum_{i=n}^{m(n,T)}\alpha_{i+1}{\bm{\epsilon}}_{i+1}||=0,$ $\mathds{P}_{G^{*}}$-a.s., where $m(n,T)=\max\{m:\sum_{i=n}^{m}\alpha_{i+1}\leq T\}$,
  • S4
@@
S1 holds by assumption; S3 holds since $({\bm{\epsilon}}_{n})$ is a martingale difference, and S4 is obvious. We now prove S2. The set $J$ is closed and $\mathscr{l}$ is continuously differentiable. Moreover, for $\mathbf{g}\in\mathring{(\Delta^{D})}$,
@@
which leads to a contradiction. Hence S2 holds, and ${\rm d}(\mathbf{g}_{n},J)\rightarrow 0$ $\mathds{P}_{G^{*}}$-a.s., as $n\rightarrow\infty$.
To conclude, we need to show that $\mathds{P}_{G^{*}}(\liminf_{n}{\rm d}(\mathbf{g}_{n},\partial\Delta^{D})=0)=0$. Since ${\rm d}(\mathbf{g}^{*},\partial\Delta^{D})>0$ and ${\rm d}(\mathbf{g}_{n},\{\mathbf{g}^{*}\}\cup\partial\Delta^{D})\rightarrow 0$ $\mathds{P}-a.s.$, then $\mathds{P}_{G^{*}}(\liminf_{n}{\rm d}(\mathbf{g}_{n},\partial\Delta^{D})=0)>0$ if and only if there exists $M$ with $\mathds{P}_{G^{*}}(M)>0$ such that ${\rm d}(\mathbf{g}_{n}(\omega),\partial\Delta^{D})\rightarrow 0$ for every $\omega\in M$, as $n\rightarrow\infty$. Let us make this assumption as a contradiction hypothesis. Define the function $\mathcal{L}(\mathbf{g})=\sum_{i=1}^{D}\log(\mathbf{g}^{*}_{i}/\mathbf{g}_{i})\mathbf{g}^{*}_{i}$. By the contradiction hypothesis, $\mathcal{L}(\mathbf{g}_{n})(\omega)\rightarrow\infty$ as $n\rightarrow\infty$, for every $\omega\in M$. Then $E_{G^{*}}({\mathcal{L}}(\mathbf{g}_{n}))\rightarrow\infty$ as $n\rightarrow+\infty$, where $E_{G^{*}}$ denote expectation with respect to $\mathds{P}_{G^{*}}$. Let $\tilde{\mathbf{g}}_{n}=\eta_{n}\mathbf{g}_{n}+(1-\eta_{n})\mathbf{g}_{n+1}$, with the random variable $\eta_{n}$ taking values in $[0,1]$, be such that
@@
$E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\right)=\alpha_{n+1}E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{h}_{i}(\mathbf{g}_{n})+{\bm{\epsilon}}_{n+1,i})\right),$
@@
$E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}_{i}}{\mathbf{g}_{n,i}}{\bm{\epsilon}}_{n+1,i}\right)=0,$